Determine the weekly gross earnings for Michelle. Her regular rate of pay is $13.50 an hour,...

  1. Determine the weekly gross earnings for Michelle. Her regular rate of pay is $13.50 an hour, with time and a half paid for all hours over 8 in a given day. Find her gross earnings for the week.

S

M

T

W

TH

F

S

Total Hours

Reg

-

8

8

6.5

8

7.5

-

OT

-

3

1.5

-

2.5

-

-


  1. Jane Doe is single and claims 2 withholding allowances. Jane has weekly earnings of $520. Her deductions are Social Security, Medicare, SDI, SWT, FWT, Union Dues, Charitable Contributions. Use the information below to determine Jane’s weekly net pay.

  • Social Security = 6.2% of gross earnings
  • Medicare = 1.45% of gross earnings
  • SDI = 1% of gross earnings
  • SWT = 3.4% of gross earnings
  • FWT = Value from using the percentage method
  • Union Dues = $15
  • Charitable Contribution = $21
Answer #1

Answer of 1st Question:

Calculation of Gross Earning for a week of Michelle:

She gets paid $13.50 per hour in regular worktime and for overtime its half of reqular pa6 i.e.$6.75 per hour

So, total reqular hour worked = 8+8+6.5+8+7.5 = 38 hours

Overtime hours = 3+1.5+2.5 = 7 hours

Therefore, Gross Earning for the Week=

=38 hours * $13.50 + 7 hours * $6.75

=$513 + $47.25

=$ 560.25

Answer of 2nd Question:

Particulars Amount ($)
Weekly Earnings 520
Social Security = 6.2% of gross earnings (32.24)
Medicare = 1.45% of gross earnings (7.54)
SDI = 1% of gross earnings (5.20)
SWT = 3.4% of gross earnings (17.68)
FWT = Value from using the percentage method($77.90*2) (155.80)
Union Dues (15)
Charitable Contribution (21)
Weekly Net Pay 265.54